OpenStudy (anonymous):

Complete the ordered pairs so that each is a solution for the given equation x=y - 3? x= 0, 2 & 3 I need to figure out Y

OpenStudy (anonymous):

to figure out y, plug in the x value given and solve for it so for 0 0=y-3 3=y now do the same for 2 and 3

OpenStudy (anonymous):

so 2=y-3 +3 +3 y=5 3=y-3 +3 +3 y=6 ?

OpenStudy (anonymous):

Right